# Spanning Training Progress: Temporal Dual-Depth Scoring (TDDS) for Enhanced Dataset Pruning

### v1

@@ 4.3 Efficient Implementation of TDDS
When computing $\mathcal{R}(\bm{x}_{n})$ with Equation 6, Equation 8, and Equation 10, storing T-epoch predicted outputs may lead to out-of-memory problems, especially with large datasets like ImageNet (containing $1,281,167$ samples across $1,000$ classes). As a solution, we suggest an efficient calculation using an exponential moving average (EMA) weighting strategy. For any training point $t\geq K$, the variance ${\mathcal{R}}_{t}(\bm{x}_{n})$ is computed over a $K$-epoch window ranging from $t-K+1$ to $t$,
@@
Input: Training dataset $\mathbb{U}$; A network $f_{\bm{\theta}}$ with weight $\bm{\theta}$; Learning rate $\eta$; Epochs $T$; Iterations $I$; Window size $K$; Decay factor $\beta$. 1:  for $t=1$ to $T$ do 2:     for $i=1$ to $I$, Sample a mini-batch $\mathbb{B}_{i}\subset\mathbb{U}$ do 3:        Record predicted probabilities $f_{\bm{\theta}_{t}}(\bm{x}_{n})$, $\bm{x}_{n}\in\mathbb{B}_{i}$ 4:        Calculate $\Delta\ell_{t}^{n}$ for each $\bm{x}_{n}\in\mathbb{B}_{i}$triangleright Defined in Equation 10 5:        Update weight $\bm{\theta}_{i+1}=\bm{\theta}_{i}-\eta\bm{G}_{t,\mathbb{B}_{i}}$ 6:     end for 7:     if $K\leq t\leq T$ then 8:        Calculate ${\mathcal{R}}_{t}(\bm{x}_{n})$ in one window, $\bm{x}_{n}\in\mathbb{U}$$\triangleright$ Defined in Equation 11 9:        Update ${\mathcal{R}}(\bm{x}_{n})$ with EMA, $\bm{x}_{n}\in\mathbb{U}$$\triangleright$ Defined in Equation 12 10:     end if 11:  end for 12:  Select Top-M $\{\mathbb{S},r\}\leftarrow\{\left(\bm{x}_{m},\bm{y}_{m}\right),{\mathcal{R}}(\bm{x}_{m})\}_{m=1}^{M}$
@@ 5 Experiments
Figure 4: Parameter analysis of range $E$ and window size $K$ on CIFAR-100 with ResNet-18. From left to right, the corresponding pruning rates are $0.3$, $0.5$, $0.7$, $0.8$, and $0.9$.
@@ 5.1 Experiment Settings
Datasets and Networks. The effectiveness of the proposed TDDS is evaluated on three popular benchmarks, i.e., CIFAR-10, CIFAR-100 [16], and ImageNet-1K [9] with three networks, i.e., ResNet-18, ResNet-34 [12], and Swin-T [18]. We also verify the cross-architecture performance on four networks including ResNet-50 [12], VGG-16 [35], MobileNet-v2 [33], EfficientNet-B0 [39], and ShuffleNet [20].
Training Hyperparameters. Theoretically, if pruning rate $p={1-\frac{|\mathbb{S}|}{|\mathbb{U}|}}$, where $0<p\leq 1$, the training process speeds up by $\frac{1}{1-p}$ times. We consider two coreset training strategies: Strategy-E prioritizes efficiency with same-epoch training on both full dataset and coreset. Strategy-P prioritizes performance by training coreset for more epochs. If $\frac{|\mathbb{S}|}{|\mathbb{U}|}=1-p$, the training process is extended to $\frac{1}{1-p}$ times of the original one. For CIFAR-10 and -100, we train ResNet-18 for $200$ epochs with a batch size of $128$. For ImageNet, ResNet-34 and Swin-T are trained for $60$ [49] and $300$ epochs [13] with batch sizes of $256$ and $1024$, respectively. SGD optimizer with momentum of $0.9$ and weight decay of 0.0005 is used for update networks. The learning rate is initialized as $0.1$ and decays with the cosine annealing scheduler. During experiments, we found that smaller batch size boosts performance at a high pruning rate (shown in Supplementary). Thus, for $p=0.8$, we halved the batch size, and for $p=0.9$, we reduced it to one-fourth. All compared methods use the same settings. With Strategy-P, for example, training a CIFAR coreset with only $50\%$ training samples takes $400$ epochs, while keeping the other hyperparameters unchanged. Considering the time-consuming training, we only conduct Strategy-P experiments at pruning rates of $0.3$, $0.5$, $0.7$.
@@ 5.3 Parameter Analysis
Estimation Range $E$ and Window Size $K$. Due to the random initialization, models adapt to prominent features (easy samples) during early training. As training progresses, models gradually acquire more abstract and complex features (hard samples) that are challenging to capture in early training. Thus, the importance of samples should be estimated in different training phases with respect to different pruning levels. As shown in Figure 4, we analyze the influence of range $E$ and window size $K$ on performance. Apparently, for a larger pruning rate, we should select samples earlier in training. For CIFAR-100, the optimal ($p,E,K$) setting is $(0.3,90,10)$, $(0.5,70,10)$, $(0.7,60,20)$, $(0.8,20,10)$, and $(0.9,10,5)$. The parameter settings of CIFAR-10 and ImageNet are reported in Supplementary.
5.4 Ablation Studies
Gradient Magnitude vs. Gradient. We use gradient magnitudes in Equation 4. Although gradients themselves appear capable of describing optimization, our experimental findings indicate that the use of gradient magnitudes yields superior performance, as depicted in Figure 6(a). We argue that gradients often prioritize samples with frequently reversing gradients, resulting in unstable training and slower convergence.
5.5 Generalization Across Architecture

### v3

@@ 4.3 Efficient Implementation of TDDS
When computing $\mathcal{R}(\bm{x}_{n})$ with Equation 6, Equation 8, and Equation 10, storing $T$-epoch predicted outputs may lead to out-of-memory problems, especially with large datasets like ImageNet (containing $1,281,167$ samples across $1,000$ classes). As a solution, we suggest an efficient calculation using an exponential moving average (EMA) weighting strategy. For any training point $t\geq K$, the variance ${\mathcal{R}}_{t}(\bm{x}_{n})$ [^3] The constant coefficient $\eta$ has been omitted. is computed over a $K$-epoch window ranging from $t-K+1$ to $t$,
@@
Input: Training dataset $\mathbb{U}$; A network $f_{\bm{\theta}}$ with weight $\bm{\theta}$; Learning rate $\eta$; Epochs $T$; Iterations $I$ per epoch; Window size $K$; Decay factor $\beta$.
1:  for $t=1$ to $T$ do
2:     for $i=1$ to $I$, Sample a mini-batch $\mathbb{B}_{i}\subset\mathbb{U}$ do
3:        Record predicted probabilities $f_{\bm{\theta}_{t}}(\bm{x}_{n})$, $\bm{x}_{n}\in\mathbb{B}_{i}$
4:        Calculate $\Delta\ell_{t}^{n}$ for each $\bm{x}_{n}\in\mathbb{B}_{i}$triangleright Defined in Equation 10
5:        Update weight $\bm{\theta}_{i+1}=\bm{\theta}_{i}-\eta\bm{G}_{t,\mathbb{B}_{i}}$
6:     end for
7:     if $K\leq t\leq T$ then
8:        Calculate ${\mathcal{R}}_{t}(\bm{x}_{n})$ in one window, $\bm{x}_{n}\in\mathbb{U}$$\triangleright$ Defined in Equation 11
9:        Update ${\mathcal{R}}(\bm{x}_{n})$ with EMA, $\bm{x}_{n}\in\mathbb{U}$$\triangleright$ Defined in Equation 12
10:     end if
11:  end for
12:  Select Top-M $\{\mathbb{S},r\}\leftarrow\{\left(\bm{x}_{m},\bm{y}_{m}\right),{\mathcal{R}}(\bm{x}_{m})\}_{m=1}^{M}$
@@ 5 Experiments
Figure 4: Parameter analysis of range $T$ and window size $K$ on CIFAR-100 with ResNet-18. From left to right, the corresponding pruning rates are $0.3$, $0.5$, $0.7$, $0.8$, and $0.9$.
@@ 5.1 Experiment Settings
Datasets and Networks. The effectiveness of the proposed TDDS is evaluated on three popular benchmarks, i.e., CIFAR-10, CIFAR-100 [20], and ImageNet-1K [9] with three networks, i.e., ResNet-18, ResNet-34 [14], and Swin-T [22]. We also verify the cross-architecture performance on four networks including ResNet-50 [14], VGG-16 [39], MobileNet-v2 [37], and ShuffleNet [24].
Training Hyperparameters. Theoretically, if pruning rate $p={1-\frac{|\mathbb{S}|}{|\mathbb{U}|}}$, where $0<p\leq 1$, the training process speeds up by $\frac{1}{1-p}$ times. We consider two coreset training strategies: Strategy-E prioritizes efficiency with same-epoch training on both full dataset and coreset. Strategy-P prioritizes performance by training coreset for more epochs. If $\frac{|\mathbb{S}|}{|\mathbb{U}|}=1-p$, the training process is extended to $\frac{1}{1-p}$ times of the original one. For CIFAR-10 and -100, we train ResNet-18 for $200$ epochs with a batch size of $128$. For ImageNet, ResNet-34 and Swin-T are trained for $60$ [52] and $300$ epochs [15] with batch sizes of $256$ and $1024$, respectively. SGD optimizer with momentum of $0.9$ and weight decay of 0.0005 is used for update networks. The learning rate is initialized as $0.1$ and decays with the cosine annealing scheduler. During experiments, we found that smaller batch size boosts performance at high pruning rates (shown in Supplementary). Thus, for $p=0.8$, we halved the batch size, and for $p=0.9$, we reduced it to one-fourth. All compared methods use the same settings. With Strategy-P, for example, training a CIFAR coreset with only $50\%$ training samples takes $400$ epochs, while keeping the other hyperparameters unchanged. Considering the time-consuming training, we only conduct Strategy-P experiments at pruning rates of $0.3$, $0.5$, $0.7$.
@@ 5.3 Parameter Analysis
Table 5: Ablation study for Dual-depth strategy. The numbers in square brackets indicate the improvement in accuracy of the method presented in the current column compared to the corresponding method in the previous column.
CIFAR-100
$p$ EL2N EL2N+outer TDDS
30% 76.89 77.47[+00.58] 77.56[+00.09]
50% 67.57 73.93[+06.36] 74.04[+00.11]
70% 36.45 67.04[+30.59] 67.78[+00.74]
80% 17.31 58.21[+40.90] 63.01[+01.80]
90% 9.10 40.36[+31.26] 54.51[+14.15]
[…]
Estimation Range $T$ and Window Size $K$. Due to the random initialization, models adapt to prominent features (easy samples) during early training. As training progresses, models gradually acquire more abstract and complex features (hard samples) that are challenging to capture in early training. Thus, the importance of samples should be estimated in different training phases with respect to different pruning levels. As shown in Figure 4, we analyze the influence of range $T$ and window size $K$ on performance. Apparently, for a larger pruning rate, we should select samples earlier in training. For CIFAR-100, the optimal ($p,T,K$) setting is $(0.3,90,10)$, $(0.5,70,10)$, $(0.7,60,20)$, $(0.8,20,10)$, and $(0.9,10,5)$. The parameter settings of CIFAR-10 and ImageNet are reported in Supplementary.
5.4 Ablation Studies
Gradient Magnitude vs Gradient. We use gradient magnitudes in Equation 4. Although gradients themselves appear capable of describing optimization, our experimental findings indicate that the use of gradient magnitudes yields superior performance, as depicted in Figure 6(a). We argue that gradients often prioritize samples with frequently reversing gradients, resulting in unstable training and slower convergence.
Dual-Depth Strategy. To investigate the effectiveness of the proposed Dual-Depth strategy, we replace the inner depth with EL2N, and Table 5 indicates that incorporating the outer depth boosts EL2N at all considered pruning rates. Nevertheless, this enhancement still falls short of our dual-depth design, particularly evident in aggressive pruning scenarios.
6 Conclusion
Acknowledgements
This research is supported in part by the National Natural Science Foundation of China under Grant 62121001, Grant 62322117, and Grant 62371365. This research is also supported by Joey Tianyi Zhou’s A*STAR SERC Central Research Fund (Use-inspired Basic Research) and Jiawei Du’s A*STAR Career Development Fund (CDF) C233312004.
Supplementary Material
Appendix A Derivation of Equation 5
The objective in Equation 4 is to minimize the Mean Squared Error (MSE) between $\bm{\mathcal{G}}_{t,\mathbb{U}}\in\mathbb{R}^{1\times N}$ and $\tilde{\bm{\mathcal{G}}}_{t,\mathbb{S}}\in\mathbb{R}^{1\times N}$,
$\mathcal{J}=\frac{1}{T}\sum_{t=1}^{T}\|\bm{\mathcal{G}}_{t,\mathbb{U}}-\tilde{\bm{\mathcal{G}}}_{t,\mathbb{S}}\|^{2}.$ (14)
Assuming we have a complete $N$-dimension orthonormal basis,
$\bm{w}_{n}^{T}\bm{w}_{m}=\delta_{nm}=\begin{cases}1,&n=m\\ 0,&\text{ else }\end{cases},$ (15)
where $\delta_{nm}$ is the kronecker delta and $n,m=1,2,\cdots N$. Given that any vector can be represented as a linear combination of the basis vectors,
$\bm{\mathcal{G}}_{t,\mathbb{U}}=\sum_{n=1}^{N}\alpha_{tn}\bm{w}_{n}.$ (16)
According to the property of orthonormal basis, we have,
$\alpha_{tn}=\bm{\mathcal{G}}^{\top}_{t,\mathbb{U}}\cdot\bm{w}_{n}$ (17)
$\bm{\mathcal{G}}_{t,\mathbb{U}}=\sum_{n=1}^{N}\left(\bm{\mathcal{G}}^{T}_{t,\mathbb{U}}\cdot\bm{w}_{n}\right)\bm{w}_{n}.$
Our goal is to find an $M$-dimension representation $\bm{\mathcal{G}}_{t,\mathbb{S}}$,
$\tilde{\bm{\mathcal{G}}}_{t,\mathbb{S}}$ $=\sum_{n=1}^{M}\alpha_{tn}\bm{w}_{n}+\sum_{n=M+1}^{N}{b}_{n}\bm{w}_{n}.$ (18)
The second term indicates bias. With Equation 17 and Equation 18, we can calculate the difference between $\bm{\mathcal{G}}_{t,\mathbb{U}}$ and $\tilde{\bm{\mathcal{G}}}_{t,\mathbb{S}}$,
$\bm{\mathcal{G}}_{t,\mathbb{U}}-\tilde{\bm{\mathcal{G}}}_{t,\mathbb{S}}=$ $\sum_{n=1}^{N}\alpha_{tn}\bm{w}_{n}-\sum_{n=1}^{M}\alpha_{tn}\bm{w}_{n}-\sum_{n=M+1}^{N}{b}_{n}\bm{w}_{n}$ (19)
$=$ $\sum_{n=M+1}^{N}\alpha_{tn}\bm{w}_{n}-{b}_{n}\bm{w}_{n}.$
After substituting Equation 19 in Equation 14, we have
$\mathcal{J}=\frac{1}{T}\sum_{t=1}^{T}\|\sum_{n=M+1}^{N}\alpha_{tn}\bm{w}_{n}-{b}_{n}\bm{w}_{n}\|^{2}.$ (20)
Taking derivative w.r.t $\alpha_{tn}$ and ${b}_{n}$ and setting to zero, we have,
${b}_{n}=\overline{\bm{\mathcal{G}}}^{\top}_{\mathbb{U}}\cdot\bm{w}_{n},$ (21)
where $\overline{\bm{\mathcal{G}}}_{\mathbb{U}}=\frac{1}{T}\sum_{t=1}^{T}\bm{\mathcal{G}}_{t,\mathbb{U}}$. Thus, Equation 20 can be reformed as
$\mathcal{J}=$ $\frac{1}{T}\sum_{t=1}^{T}\left\|\sum_{n=M+1}^{N}\left(\bm{\mathcal{G}}^{\top}_{t,\mathbb{U}}\cdot\bm{w}_{n}-\overline{\bm{\mathcal{G}}}^{\top}_{\mathbb{U}}\cdot\bm{w}_{n}\right)\bm{w}_{n}\right\|^{2}$ (22)
$=$ $\frac{1}{T}\sum_{t=1}^{T}\left\|\sum_{n=M+1}^{N}\left((\bm{\mathcal{G}}_{t,\mathbb{U}}-\overline{\bm{\mathcal{G}}}_{\mathbb{U}})^{\top}\cdot\bm{w}_{n}\right)\bm{w}_{n}\right\|^{2}$
$=$ $\frac{1}{T}\sum_{t=1}^{T}(\sum_{n=M+1}^{N}\left((\bm{\mathcal{G}}_{t,\mathbb{U}}-\overline{\bm{\mathcal{G}}}_{\mathbb{U}})^{\top}\cdot\bm{w}_{n}\right)\bm{w}_{n})^{\top}\cdot$
$(\sum_{n=M+1}^{N}\left((\bm{\mathcal{G}}_{t,\mathbb{U}}-\overline{\bm{\mathcal{G}}}_{\mathbb{U}})^{\top}\cdot\bm{w}_{n}\right)\bm{w}_{n})$
$=$ $\frac{1}{T}\sum_{t=1}^{T}\sum_{n=M+1}^{N}\sum_{m=M+1}^{N}\left(\left(\bm{\mathcal{G}}_{t,\mathbb{U}}-\overline{\bm{\mathcal{G}}}_{\mathbb{U}}\right)^{\top}\cdot\bm{w}_{n}\right)\bm{w}_{n}^{\top}\cdot$
$\bm{w}_{m}\left(\left(\bm{\mathcal{G}}_{t,\mathbb{U}}-\overline{\bm{\mathcal{G}}}_{\mathbb{U}}\right)^{\top}\cdot\bm{w}_{m}\right)$
$=$ $\frac{1}{T}\sum_{t=1}^{T}\sum_{n=M+1}^{N}\left(\bm{\mathcal{G}}_{t,\mathbb{U}}-\overline{\bm{\mathcal{G}}}_{\mathbb{U}}\right)^{T}\cdot\bm{w}_{n}\left(\bm{\mathcal{G}}_{t,\mathbb{U}}-\overline{\bm{\mathcal{G}}}_{\mathbb{U}}\right)^{\top}\cdot\bm{w}_{n}$
$=$ $\frac{1}{T}\sum_{t=1}^{T}\sum_{n=M+1}^{N}\bm{w}_{n}^{\top}\cdot\left(\bm{\mathcal{G}}_{t,\mathbb{U}}-\overline{\bm{\mathcal{G}}}_{\mathbb{U}}\right)\left(\bm{\mathcal{G}}_{t,\mathbb{U}}-\overline{\bm{\mathcal{G}}}_{\mathbb{U}}\right)^{\top}\cdot\bm{w}_{n}.$
Minimizing $\mathcal{J}$ is equivalent to reducing the variance of the pruned samples. Consequently, the goal outlined in Equation 4 effectively becomes maximizing the variance of coreset shown in Equation 5.
Figure 7: Parameter analysis of range $T$ and window size $K$ on CIFAR-10 with ResNet-18. From left to right, the corresponding pruning rates are $0.3$, $0.5$, $0.7$, $0.8$, and $0.9$.
Table 6: Accuracy results on CIFAR-10 and 100 with smaller batch size. With a smaller batch size, all compared methods are enhanced under aggressive pruning, while our superiority remains consistent.
CIFAR-10 CIFAR-100
$p$ 80% 90% 80% 90%
batch size
128128{128}128
6464{64}64
128128{128}128
3232{32}32
128128{128}128
6464{64}64
128128{128}128
3232{32}32
Random
$86.92$
±$\scriptscriptstyle\pm 0.28$
$88.87$
±$\scriptscriptstyle\pm 0.47$
$76.71$
±$\scriptscriptstyle\pm 0.15$
$83.77$
±$\scriptscriptstyle\pm 0.27$
$56.19$
±$\scriptscriptstyle\pm 1.09$
$57.79$
±$\scriptscriptstyle\pm 0.24$
$34.88$
±$\scriptscriptstyle\pm 1.74$
$46.68$
±$\scriptscriptstyle\pm 1.07$
Entropy [8]
$80.77$
±$\scriptscriptstyle\pm 0.26$
$83.49$
±$\scriptscriptstyle\pm 0.21$
$63.65$
±$\scriptscriptstyle\pm 0.62$
$72.06$
±$\scriptscriptstyle\pm 0.81$
$38.55$
±$\scriptscriptstyle\pm 1.49$
$42.86$
±$\scriptscriptstyle\pm 0.25$
$24.09$
±$\scriptscriptstyle\pm 0.47$
$29.56$
±$\scriptscriptstyle\pm 0.54$
Forgetting [44]
$61.94$
±$\scriptscriptstyle\pm 1.33$
$76.18$
±$\scriptscriptstyle\pm 3.18$
$38.95$
±$\scriptscriptstyle\pm 0.28$
$45.87$
±$\scriptscriptstyle\pm 1.87$
$38.11$
±$\scriptscriptstyle\pm 0.55$
$38.42$
±$\scriptscriptstyle\pm 1.13$
$19.88$
±$\scriptscriptstyle\pm 0.69$
$25.82$
±$\scriptscriptstyle\pm 0.52$
EL2N [32]
$59.28$
±$\scriptscriptstyle\pm 3.62$
$68.64$
±$\scriptscriptstyle\pm 3.70$
$23.54$
±$\scriptscriptstyle\pm 0.69$
$31.89$
±$\scriptscriptstyle\pm 1.51$
$14.67$
±$\scriptscriptstyle\pm 0.94$
$17.31$
±$\scriptscriptstyle\pm 0.33$
$5.54$
±$\scriptscriptstyle\pm 0.08$
$9.10$
±$\scriptscriptstyle\pm 0.69$
AUM [34]
$59.11$
±$\scriptscriptstyle\pm 3.46$
$69.60$
±$\scriptscriptstyle\pm 3.11$
$30.62$
±$\scriptscriptstyle\pm 0.29$
$34.74$
±$\scriptscriptstyle\pm 0.11$
$16.85$
±$\scriptscriptstyle\pm 0.49$
$18.43$
±$\scriptscriptstyle\pm 0.47$
$7.99$
±$\scriptscriptstyle\pm 0.17$
$9.29$
±$\scriptscriptstyle\pm 0.27$
Moderate [47]
$86.45$
±$\scriptscriptstyle\pm 0.31$
$87.76$
±$\scriptscriptstyle\pm 0.28$
$76.11$
±$\scriptscriptstyle\pm 2.25$
$83.61$
±$\scriptscriptstyle\pm 0.24$
$54.22$
±$\scriptscriptstyle\pm 0.58$
$56.52$
±$\scriptscriptstyle\pm 0.37$
$30.50$
±$\scriptscriptstyle\pm 1.21$
$41.82$
±$\scriptscriptstyle\pm 1.11$
Dyn-Unc [15]
$73.28$
±$\scriptscriptstyle\pm 0.50$
$79.76$
±$\scriptscriptstyle\pm 1.09$
$31.99$
±$\scriptscriptstyle\pm 0.74$
$37.12$
±$\scriptscriptstyle\pm 1.12$
$36.21$
±$\scriptscriptstyle\pm 0.18$
$39.19$
±$\scriptscriptstyle\pm 0.27$
$11.68$
±$\scriptscriptstyle\pm 0.08$
$15.20$
±$\scriptscriptstyle\pm 0.41$
TDDS $\mathbf{89.82}$$\scriptscriptstyle\mathbf{\pm 0.15}$ $\mathbf{91.30}$$\scriptscriptstyle\mathbf{\pm 0.25}$ $\mathbf{77.96}$$\scriptscriptstyle\mathbf{\pm 0.29}$ $\mathbf{85.46}$$\scriptscriptstyle\mathbf{\pm 0.21}$ $\mathbf{59.56}$$\scriptscriptstyle\mathbf{\pm 0.42}$ $\mathbf{63.01}$$\scriptscriptstyle\mathbf{\pm 0.12}$ $\mathbf{51.32}$$\scriptscriptstyle\mathbf{\pm 0.16}$ $\mathbf{54.51}$$\scriptscriptstyle\mathbf{\pm 0.22}$
@@ Appendix A Derivation of Equation 5
Table 8: Cross-architecture generalization performance on CIFAR-100 with ResNet-18.
ResNet-50 VGG-16 MobileNet-v2 ShuffleNet
$p$ 30%
${50\%}$
${70\%}$
30%
${50\%}$
${70\%}$
30%
${50\%}$
${70\%}$
30%
${50\%}$
${70\%}$
Random
72.0972.09{72.09}72.09
68.2768.27{68.27}68.27
61.7561.75{61.75}61.75
71.7571.75{71.75}71.75
67.5767.57{67.57}67.57
61.0361.03{61.03}61.03
70.2770.27{70.27}70.27
67.7667.76{67.76}67.76
63.0263.02{63.02}63.02
68.3168.31{68.31}68.31
65.1765.17{65.17}65.17
58.2958.29{58.29}58.29
Entropy [8]
73.0973.09{73.09}73.09
63.1263.12{63.12}63.12
47.6147.61{47.61}47.61
69.5269.52{69.52}69.52
61.1661.16{61.16}61.16
48.4248.42{48.42}48.42
67.9167.91{67.91}67.91
61.6961.69{61.69}61.69
51.7451.74{51.74}51.74
64.1264.12{64.12}64.12
56.2856.28{56.28}56.28
44.6844.68{44.68}44.68
Forgetting [44]
78.1778.17{78.17}78.17
70.6070.60{70.60}70.60
48.7448.74{48.74}48.74
73.2973.29{73.29}73.29
66.0166.01{66.01}66.01
47.8547.85{47.85}47.85
72.3772.37{72.37}72.37
68.0568.05{68.05}68.05
54.0654.06{54.06}54.06
66.9466.94{66.94}66.94
60.6460.64{60.64}60.64
40.6540.65{40.65}40.65
EL2N [32]
76.2776.27{76.27}76.27
65.8365.83{65.83}65.83
23.3523.35{23.35}23.35
72.4272.42{72.42}72.42
63.0763.07{63.07}63.07
36.4736.47{36.47}36.47
71.9671.96{71.96}71.96
63.8163.81{63.81}63.81
42.4742.47{42.47}42.47
69.2169.21{69.21}69.21
56.8256.82{56.82}56.82
29.2229.22{29.22}29.22
AUM [34]
77.3877.38{77.38}77.38
64.264.2{64.2}64.2
32.3632.36{32.36}32.36
73.6073.60{73.60}73.60
62.0162.01{62.01}62.01
30.8830.88{30.88}30.88
72.2972.29{72.29}72.29
64.3364.33{64.33}64.33
36.3536.35{36.35}36.35
66.9866.98{66.98}66.98
54.3154.31{54.31}54.31
29.2429.24{29.24}29.24
Moderate [47]
72.6772.67{72.67}72.67
68.7568.75{68.75}68.75
57.6157.61{57.61}57.61
70.170.1{70.1}70.1
65.5665.56{65.56}65.56
57.8057.80{57.80}57.80
70.0170.01{70.01}70.01
67.0367.03{67.03}67.03
60.7860.78{60.78}60.78
66.5366.53{66.53}66.53
62.5362.53{62.53}62.53
50.3350.33{50.33}50.33
Dyn-Unc [15]
73.0073.00{73.00}73.00
63.763.7{63.7}63.7
46.2646.26{46.26}46.26
68.4868.48{68.48}68.48
61.2761.27{61.27}61.27
47.2447.24{47.24}47.24
68.0168.01{68.01}68.01
62.0462.04{62.04}62.04
49.4049.40{49.40}49.40
64.5764.57{64.57}64.57
56.1456.14{56.14}56.14
42.3542.35{42.35}42.35
CCS [52]
76.9676.96{76.96}76.96
72.4372.43{72.43}72.43
64.7464.74{64.74}64.74
74.0274.02{74.02}74.02
70.1470.14{70.14}70.14
$\mathbf{64.40}$
73.0473.04{73.04}73.04
70.6370.63{70.63}70.63
66.3166.31{66.31}66.31
69.8069.80{69.80}69.80
66.7166.71{66.71}66.71
$\mathbf{61.31}$
TDDS
$\mathbf{79.53}$
$\mathbf{76.24}$
$\mathbf{66.56}$
$\mathbf{74.23}$
$\mathbf{70.66}$
64.0864.08{64.08}64.08
$\mathbf{74.23}$
$\mathbf{71.14}$
$\mathbf{66.36}$
$\mathbf{70.14}$
$\mathbf{67.14}$
61.0761.07{61.07}61.07
Figure 8: Illustration of the different types of noise used for image corruption. Here we consider Gaussian noise, random occlusion, resolution, fog, and motion blur.
Appendix B Comparision Methods
Random randomly selects partial data from the full dataset to form a coreset.
Entropy [8] is a metric of sample uncertainty. Samples with higher entropy are considered to have a greater impact on model optimization. The entropy is calculated with the predicted probabilities at the end of training.
Forgetting [44] counts how many times the forgetting happens during the training. The unforgettable samples can be removed with minimal performance drop.
EL2N [32] selects samples with larger gradient magnitudes which can be approximated by error vector scores. Only the first $10$-epoch error vector scores are averaged to evaluate samples.
AUM [34] selects samples with the highest area under the margin, which measures the probability gap between the target class and the next largest class across all training epochs. A larger AUM suggests higher importance.
Moderate [47] calculates sample-wise distance in feature space. Samples near the median are considered more important. Here, the features are generated by a pretrained model.
CCS [52] uses a variation of stratified sampling across importance scores to improve the coverage of coreset, which can be combined with other criteria. In our experiments, AUM [34] is used as the importance measurement in CCS [52].
Dyn-Unc [15] calculates the dynamic uncertainty defined as the variance of target class predicted probabilities during the training progress. The samples with larger uncertainties are more important than those with smaller uncertainties.
Note that, we use the same experimental hyperparameter settings to ensure equity for all the compared methods.
Appendix C Parameter Settings
The grid search of CIFAR-10 is shown in Figure 7. The optimal ($p,T,K$) setting is $(0.3,70,10)$, $(0.5,90,10)$, $(0.7,80,10)$, $(0.8,30,10)$, and $(0.9,10,5)$. For ImageNet-1K with ResNet-34, we set $(0.3,20,10)$, $(0.5,20,10)$, and $(0.7,30,20)$. For ImageNet-1K with Swin-T, we set $(0.25,200,10)$, $(0.3,180,10)$, $(0.4,150,10)$, and $(0.5,100,10)$.
Appendix D Small Batch Size Boosts Aggressive Pruning
In the experiments, we reveal that smaller batch size boosts coreset training, especially under aggressive pruning rates. This phenomenon is attributed to so-called Generalization Gap [18], which suggests that when the available data is extremely scare, smaller batch size can prevent overfitting by allowing more random explorations in the optimization space before converging to an optimal minimum. As reported in Table 6, smaller batch size improves the accuracy of high pruning rates for all the compared methods. Note that, regardless of the batch size, our method consistently demonstrates a significant advantage.
Appendix E Generalization Across Architectures